OpenStudy (anonymous):

Factor! 16w^2-56w+49

OpenStudy (anonymous):

16w^2-56w+49

OpenStudy (akashdeepdeb):

Use (a+b)^2 = a^2 + 2ab + b^2 IDENTITY. 16w^2 - 56w + 49 (4w)^2 - 2(4w)(7) + (7)^2 Understood? :) NOTE: DO NOT use wolfram alpha.

OpenStudy (anonymous):

Is this the answer (4w-7)(4w-7)

OpenStudy (akashdeepdeb):

It does but we should work it out ourselves. It does NOT help in the long run. :)

OpenStudy (anonymous):

Im doing it on my notebook thats why. its more easier

OpenStudy (akashdeepdeb):

strike out the MORE for correct grammar. :)

OpenStudy (akashdeepdeb):

That is the answer. :)

OpenStudy (anonymous):

I need help again the problem is a^4-1 and my answer is (a^2-1)(a^2+1). What did i do wrong?

OpenStudy (akashdeepdeb):

Factorize (a^2-1) even further. :)

OpenStudy (anonymous):

how

OpenStudy (akashdeepdeb):

The same way you factorized (a^4-1).

OpenStudy (anonymous):

1 is a prime factor right?

OpenStudy (anonymous):

i can't get 1

OpenStudy (akashdeepdeb):

See, use the identity (a+b)(a-b) = a2 - b2 Getting the algebraic Identity? :D

OpenStudy (anonymous):

(a^2-1)(a^2+2) is this right?

OpenStudy (akashdeepdeb):

We have to simplify (a^4-1) = \[(a^4-1) = (a^2-1)(a^2+1)\] [From Identity (a+b)(a-b) = a2 - b2] Now good thing is we have \(a^2-1\) which we can factorize further using the same identity. So \(a^2-1\) can be written as (a+1)(a-1). Thus, finally the ultimate factor form of \(a^4-1= (a^2+1)(a+1)(a-1)\) Understood? :)

OpenStudy (anonymous):

Now i understand

OpenStudy (akashdeepdeb):

ok